OpenStudy (anonymous):

SIMPLIFY square root of 20^4

OpenStudy (debbieg):

Do you know the rules for rational exponents? \[\Large \sqrt{20^4}= (20^4)^{1/2}\]

OpenStudy (debbieg):

or, alternatively.... \[\Large \sqrt{20^4}=\sqrt{(20^2)^2}\]

OpenStudy (debbieg):

lol while I was responding to your comment, you deleted it.... ?

OpenStudy (anonymous):

sorry wasn't sure if it made sense..

OpenStudy (debbieg):

Well, what I was going to say is that that sounds like a convuluted approach :) but I presume that "prime number and perfect square" refers to writing 20 = 5*4, so I guess that means: \[\Large \sqrt{20^4}=\sqrt{20\cdot 20\cdot 20\cdot 20}=\sqrt{20}\cdot\sqrt{20}\cdot\sqrt{20}\cdot\sqrt{20}\] and then you can write each 20 as 5*4, so you can pull each 4 out from under the radical as the square root, 2, and then you have four factors of \(\Large \sqrt{5}\).... Like I said, really convoluted. :)

OpenStudy (anonymous):

oh my goodness.. i didn't add my x in there

OpenStudy (anonymous):

it's supposed to be; the square root of 20x^4

OpenStudy (debbieg):

Oh, well that's very different indeed. lol.

OpenStudy (debbieg):

\[\Large \sqrt{20x^4}=\sqrt{20}\cdot \sqrt{x^4}=\sqrt{4\cdot 5}\cdot \sqrt{(x^2)^2}\] Then simplify each part. you can pull out the 4, but not the 5. And you can completely simplify the radical involving the power of x.

OpenStudy (anonymous):

what do you mean pull out the 4 but not 5?

OpenStudy (debbieg):

OK, to simplify a square root expression, you have to remember that you can "pull out" any perfect squares under the square root sign. Anything under there that ISNT a square doesn't get to come out and play, it has to stat where it is. :) I'll so a DIFFERENT one for you: \[\Large \sqrt{50}=\sqrt{25\cdot 2}=\sqrt{25}\cdot \sqrt{2}=\sqrt{5^2}\cdot \sqrt{2}\] Now, the 5^2 gets to come out, because the square root of a perfect square is just what you started with, right? The square root "undoes" the squaring. But the 2 has to stay where it is, can't simplify because it isn't a perfect square... \(\sqrt{2}\) is just \(\sqrt{2}\), there is nothing more that you can do with it. so: \[\Large \sqrt{5^2}\cdot \sqrt{2}=5 \sqrt{2}\] Now, can you do the same thing for \(\sqrt{20}\)?

OpenStudy (anonymous):

i still don't fully understand how to do it for 20 bc of 4 and 5 :/

OpenStudy (debbieg):

Just like I did it for 50: factor to see what you have, under the sq root. Any perfect square comes out, because it simplifies (you take the sq. root - like I did above for 25) \[\Large \sqrt{20}=\sqrt{4\cdot 5}=\sqrt{4}\cdot \sqrt{5}=\sqrt{2^2}\cdot \sqrt{5}\] So, the \(\Large \sqrt{2^2}\) will simplify, since it's a perfect square, the 2 comes out. But the \(\Large \sqrt{5}\) stays under the radical, since 5 is not a perfect square.

OpenStudy (anonymous):

\[2\sqrt{5x ^{4}}\]

OpenStudy (debbieg):

OK, you took care of the 20 part of it. Now what about the \(\Large x^4\)? That is a perfect square, since: \(\Large x^4=(x^2)^2\).

OpenStudy (anonymous):

wouldn't it just be x?

OpenStudy (debbieg):

Not quite.... think about what is SQUARED under the SQUARE ROOT. Whatever it is that is squared under the root, that's what comes out.

OpenStudy (anonymous):

so x^2?

OpenStudy (debbieg):

\[\Large z^6=(z^3)^2~~so~~\sqrt{ z^6}=z^3\] \[\Large t^{20}=(t^{10})^2~~so~~\sqrt{ t^{20}}=t^{10}\]

OpenStudy (debbieg):

RIGHT! You got it!

OpenStudy (debbieg):

So now put it all together... what's the complete simplified form?

OpenStudy (anonymous):

\[2\sqrt{5x ^{2}}\]

OpenStudy (debbieg):

Ack.... now wait a sec.... that \(x^2\) COMES OUT, remember?

OpenStudy (debbieg):

You "take the square root" so it comes OUT of the radical... out in front... just like the square root of 4 came out, so does the square root of x^4

OpenStudy (anonymous):

ohh okay! Got it!

OpenStudy (anonymous):

\[2x ^{2}\sqrt{5}\]

OpenStudy (debbieg):

\(\begin{matrix}\Huge \color{gold}{★★★★★★★★★★}\\ \Huge \color{red}{\text{Great job!}} \\ \Huge \color{red}{\text{Gold stars for you!}}\\ \Huge \begin{matrix}\color{gold}{★★★}\\\color{gold}{★★}\\\color{gold}{★★★}\end{matrix} \color{purple}{\overbrace{\left(\begin{matrix}\color{blue}{\widehat\odot \widehat\odot} \\\small\bigtriangleup \\ \Huge \color{red}\smile\end{matrix}\right)}} \Huge \begin{matrix}\color{gold}{★★★}\\\color{gold}{★★}\\\color{gold}{★★★}\end{matrix}\end{matrix}\)

OpenStudy (debbieg):

:)

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (debbieg):

You're welcome, happy to help. :)